OpenStudy (anonymous):

AB←→ and BC←→ form a right angle at point B. If A = (-3, -1) and B = (4, 4), what is the equation of BC

OpenStudy (anonymous):

1.x+3y=16 2.2x+y=12 3.-7x-5y=-48 4.7x-5y=48

OpenStudy (triciaal):

pythagorean theorem for a right triangle with legs a and b, a^2 + b^2 = c^2 c is the length opposite the right angle to find the length of one leg find the difference in the x- coordinates to find the length of the other leg find the difference in the y-coordinates slope of the line is the change in y divided by the corresponding change in x slope is (y2-y1)/(x2-x1) given the points (x1,y1) and (x2,y2) using either point on the line find the intercept write the general equation of the line for the point (x,y) rewrite to standard form Ax + By = C

OpenStudy (anonymous):

thanks for your help!!!!!!!!

OpenStudy (triciaal):

you are welcome!